# Entanglement measures for detectability

### v2

@@ Abstract
We propose new entanglement measures as the detection performance based on the hypothesis testing setting. We clarify how our measures work for detecting an entangled state by extending the quantum Sanov theorem. Our analysis covers the finite-length setting. Exploiting this entanglement measure, we present how to derive entanglement witness to detect the given entangled state by using the geometrical structure of this measure. We derive their calculation formulas for maximally correlated states, and propose their algorithms that work for general entangled states. In addition, we investigate how our algorithm works for solving the membership problem for separability.
@@ I Introduction
Entanglement is a valuable resource for quantum information processing. It is important to distinguish a given entangled state from separable states, i.e., non-entangled states. However, detecting an entanglement state has two meanings, the physical detection and the mathematical detection. The physical detection means the experimental detection of a generated entangled state. The mathematical detection means identifying whether a given density matrix has a separable form or not. Since the mathematical detection is often called the membership problem of separability, in this paper, we consider that entanglement detection means the physical one while this paper address both problems.
@@
Several practical methods to achieve entanglement detection have been proposed [1, 2, 3, 4, 5, 6, 7], but they did not discuss the optimal performance for the detection. On the other hand, to characterize the amount of entanglement, many entangled measures have been proposed. However, these proposed entangled measures have less relation with the experimental detectability of entanglement. In fact, when Vedral et al [8, 9] introduced entanglement of relative entropy, they pointed out the relation between the hypothesis testing and the relative entropy. They employed the minimum relative entropy between the given entangled state and a separable state. However, this measure expresses the performance of the discrimination between the given entangled state and the closest separable state in the sense of relative entropy. When one hypothesis is chosen to be the set of all separable states, this measure does not express the performance of this discrimination [10]. Until now, no study discussed the optimal detection efficiency for entanglement verification.
@@
In this paper, to resolve this problem, we propose a new entanglement measure that expresses the optimal performance of the discrimination between the given entangled state and the set of separable states. In this characterization, in the first step, we employ the quantum Sanov theorem [11, 12] for hypothesis testing, and consider the same minimization as entanglement of relative entropy by exchanging the first and second inputs in the relative entropy. The proposed quantity satisfies the entanglement monotone. To enhance this characterization, we derive a much simpler upper bound of the error probability in the finite-length setting of this kind of hypothesis testing, which can be considered another proof of the quantum Sanov theorem. In addition, we derive the strong converse exponent of the same problem setting as the quantum Sanov theorem. However, when we simply apply the quantum Sanov theorem, we assume that the separable states are limited to the iid states of separable states. To remove this assumption, in the second step, we derive the optimal performance of this discrimination when possible separable states are allowed to take a form of a general separable state.
@@
This entanglement measure is useful for finding an entanglement witness to detect the target entangled state. This relation is characterized by information geometry [13]. Also. we calculate the proposed entanglement measures when a state is a maximally correlated state. This class contains all pure states. For this calculation, we introduce a notable relation with coherence measures, and several calculation formulas for coherence measures.
@@
In the next step, we propose an efficient algorithm to compute the proposed entanglement measure on the joint system ${\mathcal{H}}_{A}\otimes{\mathcal{H}}_{B}$. It is quite difficult to directly calculate entanglement measures because the calculations of entanglement measures are directly related to the membership problem of separability, which is known as NP-hard [14, 15, 16, 17]. Similar to the paper [17, Section 3], we employ $\delta$-nets on the second system ${\mathcal{H}}_{B}$. In the following, we assume that the first system ${\mathcal{H}}_{A}$ has a larger dimension. That is, to optimize the choice of a separable state, we fix as many states on ${\mathcal{H}}_{B}$ as possible, and optimize the choice of states on ${\mathcal{H}}_{A}$. Then, we define an approximate value of the proposed measure. Employing the quantum Arimoto-Blahut algorithm proposed by the paper [18], we propose an algorithm for the approximate value. This algorithm is an iterative algorithm, and we show its convergence to the global minimum and its convergence speed.
@@
As the third step, we apply our algorithm to the membership problem of separability. Although it is known that this problem is NP-hard [14, 15, 16, 17], several algorithms have been proposed for this problem. As a typical method, Doherty et al. [19, 20] proposed a method to identify whether a given state is a separable state or not by considering “symmetric extensions” and “PPT-symmetric extensions” in the second system ${\mathcal{H}}_{B}$. This kind of extension is called DPS hierarchy. Later, Navascués et al. [21] evaluated its calculation complexity. DPS hierarchy was discussed in [22, (14)], [23, 24]. The paper [17] proposed another algorithm for this problem by using $\delta$-nets on the second system ${\mathcal{H}}_{B}$. The paper [17] also discussed its calculation complexity. It was shown that the calculation complexity of these preceding algorithms are exponential only for the dimension of the second system ${\mathcal{H}}_{B}$. The papers [21, 17] evaluated the exponential component for the dimension of the second system ${\mathcal{H}}_{B}$ in their algorithms. The paper [21] evaluated the exponential component for the dimension of the first system ${\mathcal{H}}_{A}$, but the paper [17] evaluated the polynomial component for the dimension of the first system ${\mathcal{H}}_{A}$. This paper discussed our calculation complexity, and showed that the calculation complexity of our algorithm is exponential only for the dimension of the second system ${\mathcal{H}}_{B}$. In addition, we evaluated the exponential component for the dimension of the second system ${\mathcal{H}}_{B}$ in our algorithm, and evaluated the polynomial component for the dimension of the first system ${\mathcal{H}}_{A}$. As the comparison with the evaluation by [21], we find that our exponential component for the dimension of the second system ${\mathcal{H}}_{B}$ is larger than that in the analysis [21], but our polynomial component for the dimension of the first system ${\mathcal{H}}_{A}$ is smaller than that in the analysis [21]. This fact shows that our algorithm has an advantage over the algorithm based on “symmetric extensions” and “PPT-symmetric extensions” when the dimension of the second system ${\mathcal{H}}_{B}$ is fixed and only the dimension of the first system ${\mathcal{H}}_{A}$ increases.
@@
Finally, we apply our algorithm to finding a separation hyperplane. When we solve a conic linear programming with a general cone, it is a key step to find a separation hyperplane for an element that does not belong to the cone [25, 26, 27, 28]. Recently, a conic linear programming with the separable cone has been applied to several problems in quantum information [29, 30]. Therefore, it is desired to find a separation hyperplane for the separable cone. In quantum information, a separation hyperplane for the separable cone is called an entanglement witness [31, 32, 33, 34, 35]. In this paper, based on our algorithm, we also propose an algorithm to find an entanglement witness, i.e., a separation hyperplane for an entangled state.
@@
The remainder of this paper is organized as follows. Section II formulates the detection of an entangled state as a problem of hypothesis testing under the iid setting. Section III addresses this problem under the non-iid setting. Section IV explains information geometrical characterization of our proposed measure and its relation to entanglement witness. Section V proposes and studies new coherence measures as the preparation of the calculation of our entanglement measures. Section VI derives several calculation formulas for our entanglement measures for maximally entangled states based on the analysis in Section V. Section VII proposes algorithms for our entanglement measures by using $\delta$-net. Section VIII applies our results to the membership problem for separability and separation hyperplane. Section IX numerically calculates our entanglement measure, and check how our method works for the membership problem. Section IV explains that our method finds an entanglement witness. Section X presents our exponential version of quantum Sanov theorem. Section XI makes conclusions and discussions. Appendix A presents concrete constructions of $\delta$-nets. Appendix B proves the theorem for the relation between our entanglement measures and our proposed coherence measure. Appendix C proves theorems for the convergence for our algorithms. Appendix D proves our exponential version of quantum Sanov theorem. Appendix E proves our exponential evaluation of error probabilities in our problem of hypothesis testing under the non-iid setting.
@@ II Detecting entangled state with iid setting
where $D(\sigma\|\rho):=\operatorname{\mathrm{Tr}}\sigma(\log\sigma-\log\rho)$. This value is different from entanglement of relative entropy $E(\rho_{AB}):=\min_{\sigma_{AB}\in{\cal S}_{AB}}D(\rho_{AB}\|\sigma_{AB})$. This fact shows that our measure $E^{*}(\rho_{AB})$ expresses the limit of the performance to distinguish the entangled state $\rho_{AB}$ from ${\cal S}_{AB}$. In contrast, when the choice of $H_{0}$ and $H_{1}$ is exchanged, $E(\rho_{AB})$ does not have the same meaning [10]. The proof of (4) is given as follows. The part $\leq$ in (4) follows from the strong converse part of quantum Stein’s lemma. The paper [11] showed the part $\geq$ in (4). However, they did not consider the case when $\epsilon$ is exponentially small. To enhance the meaning of the quantity $E^{*}(\rho_{AB})$, it is needed to evaluate $\beta_{\epsilon,n}(\rho_{AB})$ in such a case.
Table 1: List of mathematical symbols for detecting entangled state from Section II onwards.
@@
$\overline{E}^{*}(\rho_{AB})$ Normalized minimum (12)
@@
$\overline{E}_{\alpha}^{*}(\rho_{AB})$ Normalized minimum relative (13)
@@
For this aim, using the sandwich relative entropy $D_{\alpha}(\sigma\|\rho):=\frac{1}{\alpha-1}\log\operatorname{\mathrm{Tr}}(\rho^{\frac{1-\alpha}{2\alpha}}\sigma\rho^{\frac{1-\alpha}{2\alpha}})^{\alpha}$, we generalize $E^{*}(\rho_{AB})$ as
@@
As presented in Section X, we have the following theorem.
@@
Under the limit $r\to 0$, the RHS of (6) goes to $E_{1}^{*}(\rho_{AB}):=E^{*}(\rho_{AB})$. Hence, the relation (6) recovers the inequality $\geq$ in (4). When a TP-CP map $\Lambda$ preserves the set ${\cal S}_{AB}$, these quantities satisfy the monotonicity as
@@ III Detecting entangled state with non-iid setting
As shown in Appendix E, we have the following theorem.
@@
As an inequality opposite to (4), the inequality (14) implies the following inequality.
@@
While $\overline{E}^{*}(\rho_{AB})\geq\tilde{E}^{*}(\rho_{AB})$, it is not clear whether equality holds in general. When the quantity $E_{\alpha}^{*}(\rho_{AB})$ satisfies the additivity $E_{\alpha}^{*}(\rho_{AB,1}\otimes\rho_{AB,2})=E_{\alpha}^{*}(\rho_{AB,1})+E_{\alpha}^{*}(\rho_{AB,2})$, the equations ${E}^{*}(\rho_{AB})=\overline{E}^{*}(\rho_{AB})=\tilde{E}^{*}(\rho_{AB})$ and ${E}_{\alpha}^{*}(\rho_{AB})=\overline{E}_{\alpha}^{*}(\rho_{AB})$ hold. Therefore, it is crucial to clarify when the above additivity holds.
IV Information-geometrical characterization
Given a non-separable state $\rho_{AB}\notin{\cal S}_{AB}$, we discuss how to find an entanglement witness to distinguish $\rho_{AB}$ and ${\cal S}_{AB}$. For this aim, we employ the minimizer of $E^{*}(\rho_{AB})$, i.e., the state

### v3

@@ Abstract
We propose new entanglement measures as the detection performance based on the hypothesis testing setting. We clarify how our measures work for detecting an entangled state by extending the quantum Sanov theorem. Our analysis covers the finite-length setting. Exploiting this entanglement measure, we present how to derive an entanglement witness to detect the given entangled state by using the geometrical structure of this measure. We derive their calculation formulas for maximally correlated states, and propose their algorithms that work for general entangled states. In addition, we investigate how our algorithm works for solving the membership problem for separability. Further, employing this algorithm, we propose a method to find entanglement witness for a given entangled state.
Index Terms:
entanglement measure, entanglement detection, membership problem, hypothesis testing, entanglement witness
@@ I Introduction
Entanglement is a valuable resource for quantum information processing. It is important to distinguish a given entangled state from separable states, i.e., non-entangled states. However, detecting an entanglement state has two meanings, the physical detection and the mathematical detection. The physical detection means the experimental detection of a generated entangled state. Given a density matrix $\rho_{AB}$ being entangled, when a physical system is prepared as the density matrix $\rho_{AB}$, we verify whether the generated physical state is entangled. This problem is a special case of quantum hypothesis testing. In the mathematical detection, given a density matrix $\rho_{AB}$, we study whether $\rho_{AB}$ is an entangled state or a separable state. Since the mathematical detection is often called the membership problem of separability, in this paper, we consider that entanglement detection means the physical one while this paper addresses these two different problems. This paper aims to propose entangled measures, and to present that our entangled measures work for both topics.
@@
Several practical methods to achieve entanglement detection have been proposed [1, 2, 3, 4, 5, 6, 7], but they did not discuss the optimal performance for the detection. On the other hand, to characterize the amount of entanglement, many entangled measures have been proposed. However, these proposed entangled measures have less relation with the experimental detectability of entanglement. In fact, when Vedral et al [8, 9] introduced relative entropy of entanglement, they pointed out the relation between the hypothesis testing and the relative entropy. They employed the minimum relative entropy between the given entangled state and a separable state. However, this measure expresses the performance of the discrimination between the given entangled state and the closest separable state in the sense of relative entropy. When one hypothesis is chosen to be the set of all separable states, this measure does not express the performance of this discrimination [10]. Until now, no study discussed the optimal detection efficiency for entanglement verification.
@@
In this paper, to resolve this problem, we propose a new entanglement measure that expresses the optimal performance of the hypothesis testing, i.e., discrimination between the given entangled state and the set of separable states. The reason is that quantum hypothesis testing is considered as a key topic in quantum Shannon theory because it is related to various topics in quantum Shannon theory [11, 12, 13, 14, 15, 16, 18, 19, 20]. In this characterization, in the first step, we employ the quantum Sanov theorem [21, 22] for hypothesis testing, and consider the same minimization as relative entropy of entanglement by exchanging the first and second inputs in the relative entropy. The proposed quantity is an entanglement monotone. To enhance this characterization, we derive a much simpler upper bound of the error probability in the finite-length setting of this kind of hypothesis testing, which can be considered another proof of the quantum Sanov theorem. In addition, we derive the strong converse exponent of the same problem setting as the quantum Sanov theorem. However, when we simply apply the quantum Sanov theorem, we assume that the separable states are limited to the independent and identical distributed (iid) states of separable states. To remove this assumption, in the second step, we derive the optimal performance of this discrimination when possible separable states are allowed to take the form of a general separable state.
@@
This entanglement measure works for finding an entanglement witness to detect the target entangled state. This relation is characterized by information geometry [23]. Also, we calculate the proposed entanglement measures when a state is a maximally correlated state. This class contains all pure states. For this calculation, we introduce a notable relation with coherence measures, and several calculation formulas for coherence measures.
@@
In the next step, we propose an efficient algorithm to compute the proposed entanglement measure on the joint system ${\mathcal{H}}_{A}\otimes{\mathcal{H}}_{B}$. It is quite difficult to directly calculate entanglement measures because the calculations of entanglement measures are directly related to the membership problem of separability, which is known as NP-hard [24, 25, 26, 27]. Similar to the paper [27, Section 3], we employ $\delta$-nets on the second system ${\mathcal{H}}_{B}$. In the following, we assume that the first system ${\mathcal{H}}_{A}$ has a larger or equal dimension. That is, to optimize the choice of a separable state, we fix as many states on ${\mathcal{H}}_{B}$ as possible, and optimize the choice of states on ${\mathcal{H}}_{A}$. Then, we define an approximate value of the proposed measure. Employing the quantum Arimoto-Blahut algorithm proposed by the paper [28], we propose an algorithm for the approximate value. This algorithm is an iterative algorithm, and we show its convergence to the global minimum and its convergence speed.
@@
As the third step, we apply our algorithm to the membership problem of separability. Although it is known that this problem is NP-hard [24, 25, 26, 27], several algorithms have been proposed for this problem. As a typical method, Doherty et al. [29, 30] proposed a method to identify whether a given state is a separable state or not by considering “symmetric extensions” and “PPT-symmetric extensions” in the second system ${\mathcal{H}}_{B}$. This kind of extension is called DPS hierarchy. Later, Navascués et al. [31] evaluated its calculation complexity. DPS hierarchy was discussed in [32, (14)], [33, 34]. The paper [27] proposed another algorithm for this problem by using $\delta$-nets on the second system ${\mathcal{H}}_{B}$. The paper [27] also discussed its calculation complexity. It was shown that the calculation complexity of these preceding algorithms are exponential only for the dimension of the second system ${\mathcal{H}}_{B}$. The papers [31, 27] evaluated the exponential component for the dimension of the second system ${\mathcal{H}}_{B}$ in their algorithms. The paper [31, Eq. (38)] evaluated the polynomial component for the dimension of the first system ${\mathcal{H}}_{A}$. This paper discussed our calculation complexity, and showed that the calculation complexity of our algorithm is exponential only for the dimension of the second system ${\mathcal{H}}_{B}$. In addition, we evaluated the exponential component for the dimension of the second system ${\mathcal{H}}_{B}$ in our algorithm, and evaluated the polynomial component for the dimension of the first system ${\mathcal{H}}_{A}$. As the comparison with the evaluation by [31], we find that our exponential component for the dimension of the second system ${\mathcal{H}}_{B}$ is larger than that in the analysis [31], but our polynomial component for the dimension of the first system ${\mathcal{H}}_{A}$ is smaller than that in the analysis [31]. This fact shows that our algorithm has an advantage over the algorithm based on “symmetric extensions” and “PPT-symmetric extensions” when the dimension of the second system ${\mathcal{H}}_{B}$ is fixed and only the dimension of the first system ${\mathcal{H}}_{A}$ increases.
@@
Finally, we apply our algorithm to finding a separation hyperplane. When we solve a conic linear programming with a general cone, it is a key step to find a separation hyperplane for an element that does not belong to the cone [35, 36, 37, 38]. Recently, a conic linear programming with the separable cone has been applied to several problems in quantum information [39, 40]. Therefore, it is desired to find a separation hyperplane for the separable cone. In quantum information, a separation hyperplane for the separable cone is called an entanglement witness [41, 42, 43, 44, 45]. From the viewpoint of entanglement theory, it is strongly desired to establish a method to find an entanglement witness for a given state, but such a method has been established only for the limited cases [46, 47]. In this paper, based on our algorithm, we also propose an algorithm to find an entanglement witness, i.e., a separation hyperplane for an entangled state. Although the method based on DPS hierarchy does not directly yield an entanglement witness, our method directly directly yields it, which is an advantage of our method.
Overall, in addition to the proposal of the new entangled measures, the contributions of this paper are summarized as follows.
[Statistical hypothesis testing] We have shown that the proposed measures present the performance of physical detection in the sense of hypothesis testing, which has not been pointed out before. (Sections II and III)
[Exponential quantum Sanov theorem] To support the above fact, we have derived the exponential version of the quantum Sanov theorem [21, 22]. This extended version clarifies the relation between the hypothesis testing with composite hypothesis and the minimum sandwiched relative Rényi entropy, which was not discussed in the existing studies [21, 22]. Thanks to this improved evaluation, we could derive the performance of the hypothesis testing when $H_{0}$ hypothesis is the set of all separable states over $n$ bipartite systems, as opposed to an IID $n$-fold tensor product of a separable state. (Sections X and III)
[Maximally correlated states] We have calculated the proposed measures when the entangled state is maximally correlated. For this calculation, we have calculated the similar value for the coherence. (Sections V and VI)
[Algorithm & membership problem] We have shown that the proposed measures can be calculated by using the generalized Arimoto-Blahut algorithm due to the good shape of the proposed measure, which has not been pointed out before. This fact shows that the proposed measures can be used for solving the membership problem for separable states. The superiority between this method and the existing method [29, 30, 31] for the membership problem depends on the parameters of this problem. In addition, we have evaluated the relation between the calculation complexity and the precision of one of the proposed algorithms. We have made a numerical demonstration. (Sections VII, VIII, and IX)
[Finding entanglement witness] We have revealed the geometrical structure of the proposed measure by using the viewpoint of information geometry. Using this geometrical structure, we have proposed a method to find an entanglement witness by using the proposed measure. This fact shows that the calculation of the proposed measure can be used for finding an entanglement witness of a given state. Applying the algorithm stated in (A4), we have proposed an algorithm to find an entanglement witness and such a method has been strongly desired from the viewpoint of entanglement theory [46, 47]. In addition, information-geometric characterization enables us to simplify the proof of the additivity of the proposed measure while the additivity was shown by [55]. (Sections IV and VIII)
@@
The remainder of this paper is organized as follows. Section II formulates the detection of an entangled state as a problem of hypothesis testing under the iid setting. Section III addresses this problem under the non-iid setting. Section IV explains information geometrical characterization of our proposed measure and its relation to an entanglement witness. Section V proposes and studies new coherence measures as the preparation of the calculation of our entanglement measures. Section VI derives several calculation formulas for our entanglement measures for maximally entangled states based on the analysis in Section V. Section VII proposes algorithms for our entanglement measures by using $\delta$-net. Section VIII applies our results to the membership problem for separability and separation hyperplane. Section IX numerically calculates our entanglement measure, and check how our method works for the membership problem. Section IV explains that our method finds an entanglement witness. Section X presents our exponential version of quantum Sanov theorem. Section XI makes conclusions and discussions. Appendix A presents concrete constructions of $\delta$-nets. Appendix B proves the theorem for the relation between our entanglement measures and our proposed coherence measure. Appendix C proves theorems for the convergence for our algorithms. Appendix D proves our exponential version of quantum Sanov theorem. Appendix E proves our exponential evaluation of error probabilities in our problem of hypothesis testing under the non-iid setting.
@@ II Detecting entangled state with iid setting
where $D(\sigma\|\rho):=\operatorname{\mathrm{Tr}}\sigma(\log\sigma-\log\rho)$. This value is different from relative entropy of entanglement $E(\rho_{AB}):=\min_{\sigma_{AB}\in{\cal S}_{AB}}D(\rho_{AB}\|\sigma_{AB})$. This fact shows that our measure $E^{*}(\rho_{AB})$ expresses the limit of the performance to distinguish the entangled state $\rho_{AB}$ from ${\cal S}_{AB}$. In contrast, when the choice of $H_{0}$ and $H_{1}$ is exchanged, $E(\rho_{AB})$ does not have the same meaning [10]. The proof of (4) is given as follows. The part $\leq$ in (4) follows from the strong converse part of quantum Stein’s lemma. The paper [21] showed the part $\geq$ in (4). However, they did not consider the case when $\epsilon$ is exponentially small. To enhance the meaning of the quantity $E^{*}(\rho_{AB})$, it is needed to evaluate $\beta_{\epsilon,n}(\rho_{AB})$ in such a case.
TABLE I: List of mathematical symbols for detecting entangled state from Section II onwards.
@@
$\overline{E}^{*}(\rho_{AB})$ Regularized minimum (12)
@@
$\overline{E}_{\alpha}^{*}(\rho_{AB})$ Regularized minimum relative (13)
@@
For this aim, using the sandwiched relative Rényi entropy $D_{\alpha}(\sigma\|\rho):=\frac{1}{\alpha-1}\log\operatorname{\mathrm{Tr}}(\rho^{\frac{1-\alpha}{2\alpha}}\sigma\rho^{\frac{1-\alpha}{2\alpha}})^{\alpha}$, we generalize $E^{*}(\rho_{AB})$ as
@@
As presented in Section X, we have the following strengthened version of (4).
@@
Inequality (6) presents an attainable exponential rate of the failure probability for detecting the entangled state $\rho_{AB}$ when we impose an exponential constraint for the incorrect detection of any separable state under the iid condition. Under the limit $r\to 0$, the RHS of (6) goes to $E_{1}^{*}(\rho_{AB}):=E^{*}(\rho_{AB})$. Hence, the relation (6) recovers the inequality $\geq$ in (4). Inequality (7) presents an upper bound for the exponential rate of the failure probability to detect the entangled state $\rho_{AB}$ when we relax the constraints for the incorrect detection of any separable state under the iid condition. Due to this inequality, when the entangled state $\rho_{AB}$ can detected with a failure probability of a strictly larger exponential rate than $e^{-nE^{*}(\rho_{AB})}$, a separable state is incorrectly detected as an entangled state with a high probability.
When a TP-CP map $\Lambda$ preserves the set ${\cal S}_{AB}$, these quantities satisfy the monotonicity as
@@
Our new measure $E^{*}(\rho_{AB})$ has a completely different behavior from conventional measures of entanglement. Most of conventional measures of entanglement of a pure state $\rho_{AB}$ take the von Neumann entropy of its reduced density. However, $E^{*}(\rho_{AB})$ becomes infinity when $\rho_{AB}$ is a pure entangled state because the support of no separable state is included in the support of $\rho_{AB}$. Although the log-negativity of a pure state $\rho_{AB}$ does not equal the von Neumann entropy of its reduced density, it takes a finite value.
Remark 1
When $H_{0}$ and $H_{1}$ are exchanged, $E(\rho_{AB})$ does not have the same meaning as $E^{*}(\rho_{AB})$. This is because quantum Sanov theorem works only when $H_{1}$ is composed of several states [10]. That is, the choice of the hypothesis composed of several states is crucial in this scenario. Instead, to prove that $E(\rho_{AB})$ upper bounds the entanglement distillation rate, the paper [54] employs the simple hypotheses that $H_{0}$ is $\sigma_{AB}^{\otimes n}$ and $H_{1}$ is $\rho_{AB}^{\otimes n}$, where $\sigma_{AB}$ is an arbitrary chosen element of ${\cal S}_{AB}$. More precisely, an application of the strong converse of Stein’s lemma to this hypothesis testing yields the strong converse for the entanglement distillation rate [51, Proof of Theorem 8.7].
@@ III Detecting entangled state with non-iid setting
As shown in Appendix E, we have the following theorem, which is an enhanced version of Theorem 1, which strengthens (13).
@@
As an extended version of (6), Inequality (14) presents an attainable exponential rate of the failure probability for detecting the entangled state $\rho_{AB}$ when we impose an exponential constraint for the incorrect detection of any separable state without the iid condition. Under the limit $r\to 0$, the RHS of (14) goes to $\lim_{\delta\to+0}\overline{E}_{1-\delta}^{*}(\rho_{AB})$. Inequality (14) implies the following inequality.
@@
While $\overline{E}^{*}(\rho_{AB})\geq\tilde{E}^{*}(\rho_{AB})$, it is not clear whether equality holds in general. The reference [55, Proposition 2] showed that the additivity $E^{*}(\rho_{AB,1}\otimes\rho_{AB,2})=E^{*}(\rho_{AB,1})+E^{*}(\rho_{AB,2})$, which implies ${E}^{*}(\rho_{AB})=\overline{E}^{*}(\rho_{AB})$ while its simple proof is given in Section IV-B.
When the quantity $E_{\alpha}^{*}(\rho_{AB})$ satisfies the additivity
$E_{\alpha}^{*}(\rho_{AB,1}\otimes\rho_{AB,2})=E_{\alpha}^{*}(\rho_{AB,1})+E_{\alpha}^{*}(\rho_{AB,2})$ (19)
for $\alpha<1$, the equations ${E}^{*}(\rho_{AB})=\overline{E}^{*}(\rho_{AB})=\tilde{E}^{*}(\rho_{AB})$ and ${E}_{\alpha}^{*}(\rho_{AB})=\overline{E}_{\alpha}^{*}(\rho_{AB})$ hold. Therefore, it is crucial to clarify when the above additivity holds. For example, as shown in (63) of Section VI, the maximally correlated state satisfies the above additivity. The reference [56, Section 5] showed that the additivity (19) with $\alpha\in[1/2,1)$ does not hold in general by considering the bipartite antisymmetric (Werner) state [57, 58] [^1] Since our sandwiched relative entropy of order $\alpha^{\prime}$ corresponds to $\tilde{D}_{1-\alpha^{\prime}}$ of [56, (10)], i.e., the case with $(\alpha,z)=(1-\alpha^{\prime},\alpha^{\prime})$. Hence, the case with $21>\alpha^{\prime}\geq 1/$ 2 belongs to the set ${\cal D}$ defined in [56, Section 2]. Hence, the discussion [56, Section 5] covers the violation of (19) the additivity (19) with $\alpha\in[1/2,1)$..
Inequality (15) presents an upper bound for the exponential rate of the failure probability to detect the entangled state $\rho_{AB}$ when we relax the constraints for the incorrect detection of any separable state without the iid condition.
IV Information-geometrical characterization: Finding entanglement witness
IV-A Single system
Our proposed measure is useful not only for hypothesis testing of the above hypotheses, but also for finding entanglement witness. That is, given a non-separable state $\rho_{AB}\notin{\cal S}_{AB}$, it is important to find an entanglement witness to distinguish $\rho_{AB}$ and ${\cal S}_{AB}$. To see this merit, we discuss its information-geometrical characterization, which yields entanglement witness.
The papers [46, 47] proposed methods to find an entanglement witness for special states. As another method, the references [59], [60, Section 16.3.3] proposed a method to use the tangent space of the set of separable states based on the Hilbert Schmidt norm $\|\cdot\|_{HS}$. The key point of the above method is the fact that the Hilbert-Schmidt norm forms the Euclidean geometry over the space of Hermitian matrices. The above method needs to find the closest separable state in the sense of the Hilbert Schmidt norm $\|\cdot\|_{HS}$. However, no method efficiently minimizes the Hilbert Schmidt norm $\|\cdot\|_{HS}$ between a given state and the set of separable states, i.e., efficiently finds
$\mathop{\rm argmin}\limits_{\sigma_{AB}\in{\cal S}_{AB}}\|\sigma_{AB}-\rho_{AB}\|_{HS}$ (20)
when the system dimensions are large. Hence, it is not easy to find an entanglement witness by using the Hilbert-Schmidt norm under large Hilbert spaces.
Fortunately, this section shows how the minimization $\min_{\sigma_{AB}\in{\cal S}_{AB}}D(\sigma_{AB}\|\rho_{AB})$ works for finding entanglement witness by employing the minimizer of $E^{*}(\rho_{AB})$, i.e., the state